The true discount on a bill is 3/4 of the banker’s discount. If the rate of interest is 10% (simple), find the time.
1
Expert's answer
2020-06-15T17:45:05-0400

"{BD\over TD}=1+(r×t)"

"{{3\over4}\over{1\over4}}=1+({10\over100}×t)"

"{0.75\over0.25}=1+0.1t"

"3=1+0.1t"

"t=20"